# supplement

### v1

@@ Abstract
Figs. 4e-g show a cavity that operates in a "convex lens" regime, focusing polaritons at a point distant from the cavity, similar to the triangular cavities. One can observe a focal point where the curvature of the wavefront changes sign from positive to negative as it passes through the focal point. Fig. 4h illustrates a schematics of the "convex lens"-like focusing. The position of the focal point and the confinement may be adjusted by the size of the cavity. For a 0.6-$\mu$m-radius cavity, the confinement reaches $\lambda/55$ (see Extended Data Fig. S8). Furthermore, several cavities launching criss-crossing PhPs can be used for the formation of the interference patterns (see Extended Data Fig. S9) forming a lattice of $(\textrm{Re E}_{\textrm{z}})^{2}$ minima and maxima with a period equal to the polariton wavelength at the intersection. Thus, using the interference of cavity-launched polaritons with various wavefronts, it is possible to form a variety of polaritonic lattices for structured light-matter interaction and opto-mechanical purposes.
@@
Discussion. To accelerate the technological application of spatially defined PhP modes, it is necessary to find a way to precisely control their spatial distribution and focus them at specific real-space locations. Some substantial progress has been made in tuning and focusing of the PhPs using in-plane anisotropic hyperbolic materials such as MoO341, 42, 43. In the case of the uniaxial hyberbolic hBN crystals, most efforts have involved a use of substrates with an in-plane anisotropy in the relevant spectral ranges 44, 45, 46. However, the influence of these substrates is generally limited. An elegant solution has been proposed by Chaudhary et. al. 30 by using phase-fluid materials (Ge3Sb2Te6 on a CaF2 substrate) that locally change their dielectric constants and thus facilitate polaritonic waveguides and metalenses. However, the focusing effect remained rather hindered and the scalable application is somewhat elusive.
@@
In conclusion, we have developed hBN sub-wavelength resonators that are compatible with standard lithographic techniques, enabling the launch of focused, long-lived PhPs. We fabricate the subwavelength hBN cavities on the interfaces between Au and SiO2 partially transparent for PhPs. The localized modes showed a significant increase of the electric field and launched polaritons that remain confined in the real-space defined by the shapes and sizes of the cavities. Using s-SNOM, we studied the real-space optical contrast of the cavity-launched PhPs. Based on the PhP’s wavelength, we demonstrated coupling to the far-field component of the mid-IR light. The far-field component, which is resonantly enhanced in the cavities, gains enough momentum to launch polaritons into the bulk of the hBN/Au heterostructures through the scattering process at the SiO2/Au interfaces. The shapes of our resonators determine the focal point location as well as the wavefront shapes. By creating different forms of nano-resonators, we showed a focusing effect similar to a "concave lens" and a "convex lens." In particular, the "concave lens"-like cavities enable field confinement down to $\lambda/55$ at the focal points. By this, we realize a novel technique for generating and focusing spatially confined easy-plane PhPs compatible with the standard silicon technology.
@@ Methods
hBN flakes were prepared via mechanical exfoliation using PDMS adhesive tapes and were transferred on Si/SiO2 substrates for selection. Relying on optical contrast, we selected flakes of suitable thicknesses and transferred them on the pre-patterned Si++/SiO2 substrates using the PMDS/PC dry transfer method47, 48.
@@
We used a commercially available scattering-type scanning near-field microscope (s-SNOM) developed by Neaspec/Attocube (cryo-neaSCOPE) with a pseudo-heterodyne mode to a detect a background-free optical response ($\Omega\pm NM$) using the decoupled optical amplitude and phase49, 50, 51. PtIr-coated AFM tips (Nanoworld, 23 nm coating) with a resonant frequency $\approx$ 250 kHz were used to obtain the topography and the optical maps. AFM tips were illuminated by a tunable commercially available mid-infrared quantum cascade laser (QCL) using 1-2 mW of power (MIRcat, $\lambda$ = 5.5-11.5 $\mu$m from the DRS Daylight Solutions Inc.) All optical measurements presented in the manuscript were obtained on the 3rd or the 4th harmonics (i.e., S3 and S4, respectively) to ensure the background-free detection.
@@
To describe the optical phenomena in PhPs at the hBN/SiO2 and hBN/Au interfaces, we employ numerical calculations in COMSOL Multiphysics, Wave Optics module. The dispersions of the propagation and attenuation constants of the PhP modes alongside the complex effective refractive indices of these modes are calculated in a 3-layer structure, where the top layer is presented by air, the second layer is presented by hBN which is an anisotropic (along the hard axis) material with permittivity tensor taken from 52 and optical axis directed perpendicular to the layers, and the third layer is either SiO2 or Au with refractive indices taken from 53, 54. The obtained effective refractive indices are used as refractive indices for hBN/SiO2 and hBN/Au domains in the 2D model, and the surrounding space is treated as a dielectric with a refractive index of 1. As a source for the 2D model, we specify the surface current density at the edges of the hBN/SiO2, hBN/Au domains, and the surrounding space.

### v2

@@ Abstract
Using the same approach, we developed a "concave lens"-like cavity (see Extended Data Fig. S9). The negative curvature of the reflecting plane produces a divergent PhP wavefront and defocusing. Furthermore, multiple cavities launching crisscrossing PhPs can be used for the formation of interference patterns (see Extended Data Fig. S10), forming a lattice of $(\textrm{Re},E_{\textrm{z}})^{2}$ minima and maxima with a period equal to the PhP wavelength at the intersection. Thus, using the interference of cavity-launched PhP with adjusted wavefronts, it is possible to engineer polaritonic lattices for structured light-matter interaction.
@@
Discussion. To accelerate the technological application of spatially defined PhP modes, it is crucial to develop a method for precisely controlling their spatial distribution and focusing them at specific real-space locations. Significant progress has been made in tuning and focusing PhPs using in-plane-anisotropic hyperbolic materials such as MoO3 46, 47, 31. In the case of uniaxial hyperbolic hBN crystals, most efforts have relied on substrates with in-plane anisotropy in the relevant spectral ranges 48, 49, 50.
A particularly elegant solution was proposed by Chaudhary et al. 30, employing a phase-change material (Ge3Sb2Te6 on a CaF2 substrate) to locally modify dielectric constants and focus tip-launched PhPs. However, the influence of such substrates is generally limited, and integrating such uncommon materials into standard silicon-based technology remains challenging. Consequently, the focusing effect remains constrained, and scalable applications are elusive.
One of the most widely adopted methods for focusing PhPs and coupling them to the far-field component of light is the use of metallic scatterers, typically made of gold 51. This approach is essential because the tip must not be present in operational devices. By employing scatterers with curved surfaces, PhPs can be focused via wave interference. However, this approach has a fundamental limitation: the strength of light coupling to PhP modes depends heavily on the angle between light polarization and the normal to the scatterer surface at the launch point. Consequently, curved scatterers can only efficiently focus PhP modes within a narrow angular range 51, making them unsuitable for achieving arbitrary PhP lensing in isotropic materials like hBN.
Our approach overcomes these limitations by utilizing standard, technologically accessible materials (Si/SiO2, Ti/Au) as a platform for PhP engineering. We introduce hBN subwavelength cavities at the Au/SiO2 interface, which enable the generation of strongly focused (down to $\lambda/70$) and spatially confined PhPs. Crucially, our approach resolves the "scatterer" issue mentioned earlier: the Au/SiO2 interface facilitates coupling to far-field light but does not actively participate in the focusing process itself. Instead, the focusing mechanism is governed by the selection of PhPs with specific $k$-vector distributions inside the cavity. As a result, for a fixed Au/SiO2 interface orientation, cavities of varying shapes and sizes can produce different focusing effects (e.g., convex, concave, or straight wavefronts) and different angular ranges of collection. Thus, we believe our approach addresses several persistent challenges in PhP photonics and paves the way for numerous future discoveries.
@@
In conclusion, we have developed hBN sub-wavelength cavities that are compatible with standard lithographic techniques, enabling the launch of focused, long-lived PhPs. We fabricate the subwavelength hBN cavities at the interfaces between Au and SiO2 partially transparent for PhPs. The localized modes exhibit a significant enhancement of the electric field inside the cavities, launching PhPs that remain spatially confined, with their confinement determined by the shape and size of the cavities. Using s-SNOM, we studied the real-space optical contrast of the cavity-launched PhPs. Based on the PhP’s wavelength, we demonstrated coupling to the far-field component of the mid-IR light. The far-field component, which is resonantly enhanced in the cavities, acquires sufficient momentum to launch polaritons into the bulk of the hBN/Au heterostructures through the scattering process at the SiO2/Au interfaces. The shape of our resonators determines both the focal point location and the wavefront geometry. By creating different forms of cavities, we showed a focusing effect similar to a "concave lens" and a "convex lens." In particular, the "convex lens"-like cavities enable field confinement down to $\lambda/70$ at the focal point. By this, we realize a novel technique for generating and focusing spatially confined in-plane PhPs compatible with the standard fabrication approaches.
@@ Methods
hBN flakes were prepared via mechanical exfoliation using PDMS adhesive tapes and were transferred on Si/SiO2 substrates for selection. Relying on optical contrast, we selected flakes of suitable thicknesses and transferred them on the pre-patterned Si++/SiO2 substrates using the PMDS/PC dry transfer method52, 53.
@@
We used a commercially available scattering-type scanning near-field microscope (s-SNOM) developed by Neaspec/Attocube (cryo-neaSCOPE) with a pseudo-heterodyne mode to a detect a background-free optical response ($\Omega\pm NM$) using the decoupled optical amplitude and phase54, 55, 56. PtIr-coated AFM tips (Nanoworld, 23 nm coating) with a resonant frequency $\approx$ 250 kHz were used to obtain the topography and the optical maps. AFM tips were illuminated by a tunable commercially available mid-infrared quantum cascade laser (QCL) using 1-2 mW of power (MIRcat, $\lambda$ = 5.5-11.5 $\mu$m from the DRS Daylight Solutions Inc.) All optical measurements presented in the manuscript were obtained on the 3rd or the 4th harmonics (i.e., S3 and S4, respectively) to ensure the background-free detection.
@@
To describe the optical phenomena in PhPs at the hBN/SiO2 and hBN/Au interfaces, we employ numerical calculations in COMSOL Multiphysics, Wave Optics module. The dispersions of the propagation and attenuation constants of the PhP modes alongside the complex effective refractive indices of these modes are calculated in a 4-layer structure, where the top layer is presented by air, the second layer is presented by hBN which is an anisotropic (along the hard axis) material with permittivity tensor taken from 57 and optical axis directed perpendicular to the layers, the third layer is a narrow air gap, and the fourth layer is either SiO2 or Au with refractive indices taken from 58, 59. The obtained effective refractive indices are used as refractive indices for hBN/SiO2 and hBN/Au domains in the 2D model, and the surrounding space is treated as a dielectric with a refractive index of 1. As a source for the 2D model, we specify the surface current density at the edges of the hBN/SiO2, hBN/Au domains, and the surrounding space.